OpenStudy (anonymous):

what is the derivative: t^3-4t^2+5t(square t) -5 I can get 3t^2-8t (not sure what to do with 5t (square t)

OpenStudy (anonymous):

by (square t) do you mean t^2?

OpenStudy (anonymous):

2t\[\sqrt{t}\] I can't type them next to one another

OpenStudy (anonymous):

it is 2t (square t)

OpenStudy (amistre64):

5t sqrt(t) = 5t t^(1/2) = 5t^(3/2)

OpenStudy (amistre64):

5*(3/2) t^((3-2)/2) = (15/2) t^(1/2) = 15sqrt(t)/2

OpenStudy (amistre64):

it tends to be easier to work with "exponents" than with radicals.... so turn all the radicals into the appropriate "exponent" and solve from there :)

OpenStudy (anonymous):

thanks